# A Joint Learning Model with Variational Interaction for Multilingual Program Translation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: IEEE/ACM International Conference on Automated Software Engineering; Oct.27-Nov.1, 2024; Sacramento, California, USccs: Software and its engineering Automatic programming
@@ 1. Introduction
Programs form the foundation of computer applications. Various programming languages have been invented to address diverse requirements. Developing a new application from scratch is time- and labor-intensive, whereas referring to and combining existing code is more efficient. Unfortunately, when combining programs written in different programming languages, the developer usually suffers from the intensive labor of manually translating programs from one programming language to another. For example, many industries spend hundreds of millions of dollars to convert code written in older programming languages (e.g., FORTRAN and COBOL) to newer ones (e.g., Java, C++) (Rozière et al., 2020). To alleviate the burden of program migration and facilitate the development of software systems, program translation, which aims to automatically translate the program from one programming language to another, has drawn significant attention in the software mining community (Ledley and Wilson, 1962; Kennedy III, 1987; Ahmad et al., 2023; Zhu et al., 2023; Yan et al., 2023; Weisz et al., 2022; Rozière et al., 2020; Zhu et al., 2022; Waters, 1988; Chu, 1993).
@@
In this paper, we argue that jointly training a unified model to translate code across multiple programming languages is superior to separately training with pairwise data. A program implemented in different programming languages should exhibit the invariant language-shared semantics, i.e., perform the same tasks. Thus, joint training on multiple different implementations of a program can complement and enhance the learning of the language-shared representation. By constructing the unified language-shared latent space, the rich resource languages can benefit the low resource languages in the translation. However, joint learning involves a new challenge: as shown in Fig. 1, practical datasets always is semi-parallel, which contain only a few complete multi-parallel samples that include all languages, while most are partially missing implementations in some programming languages. To tackle the issue of semi-parallel data, utilizing semi-supervised learning techniques is intuitive. Generative semi-supervised learning offers a solution: all data, including partially missing and multi-parallel data, can be generated from a latent distribution. All these data can be used to learn the latent distribution and generative process with an Expectation–Maximization (EM)-based algorithm (Kingma et al., 2014). Once the latent distribution and generative process are learned, program translation can be achieved with a conditional generative process.
@@
    We conducted extensive experiments to demonstrate the effectiveness of VIM-PT and the necessity of jointly learning a unified model in multilingual program translation.
@@ 2.1. Program Translation
To meet some specific requirements, rule-based translation methods have been developed (Alsubhi et al., 2019; An et al., 2018; Pino et al., 2017; Tauda et al., 2021). Further, some works applied phrase-based statistical machine translation techniques to program translation (Karaivanov et al., 2014; Nguyen et al., 2015, 2013, 2016; Allamanis et al., 2018; Oda et al., 2015b, c, a), which leveraged grammatical structures of programming languages for code migration. Recently, various deep learning techniques were employed to program translations (Lample et al., 2018; Wen et al., 2022; Weisz et al., 2021), which can be distinguished into pairwise program translation methods (Chen et al., 2018; Guo et al., 2021; Du et al., 2023) and multilingual program translation methods (Zhu et al., 2022) by using bilingual or multilingual corpus in the supervised learning.
The pairwise approaches trained one model for each translation direction independently (Guo et al., 2021; Sun et al., 2022; Liu et al., 2024). However, a problem encountered by many existing models is that program translation datasets are usually not balanced in size for all the languages. Some languages may have much less parallel data than others. Less parallel training data can significantly affect the translation performance of low-resource languages. Therefore, some works leveraged multilingual training to improve the performance of low-resource languages, which is defined as multilingual program translation. However, the existing multilingual program translation model MuST-PT (Zhu et al., 2022) is limited to the pairwise training process and does not take full advantage of the multi-parallel data and partially missing data (e.g., non-parallel data).
@@ 2.2. Exploring Non-Parallel Data for Translation
The quality of translation systems highly depends on the quality of the available parallel data. However, for most languages, parallel resources are rare or nonexistent. Some approaches have investigated the use of non-parallel data to improve existing program translation systems (Rozière et al., 2020, 2022; Szafraniec et al., 2022; Lachaux et al., 2021; Feng et al., 2020; Wang et al., 2021). For instance, Lachaux et al. (2021) proposed DOBF that leverages the structural aspect of coding languages and pre-trains a model to recover the original version of obfuscated source code. Feng et al. (2020) proposed CodeBERT, which is a Bert-like model pre-trained in an open-source GitHub repository and used in many down-stream tasks, and then much relevant research emerged (Guo et al., 2021; Du and Yu, 2023). Rozière et al. (2020) proposed TransCoder, which trains a translation model only using the monolingual corpus with back-translation and denoising auto-encoding.
@@ 3.1. Generative Model via Variational Inference
Consider a sample program implemented in $N$ languages, denoted as $(x^{1},x^{2},\cdots,x^{N})\sim p_{D}(x^{1},x^{2},\cdots,x^{N})$, where $x_{i}$ represents the code written in the $i$-th programming language. All of these $N$ codes exhibit the same semantics, i.e., perform the same task, while each code $x_{i}$ involves the language-specific grammatical style. From the perspective of disentanglement, each code is generated through the interplay of a language-shared feature, denoted as $z^{s}\in Z^{s}$, and a language-specific feature, denoted as $z^{i}\in Z^{i}$ for the $i$-th programming language.
@@ 4.1. Dataset Description
We conduct experiments on the CoST dataset (Zhu et al., 2022), which is a large and comprehensive dataset to evaluate the performance of program translation approaches. The dataset consists of both snippet-level and program-level parallel data from 7 programming languages (i.e., C, C#, C++, Java, JavaScript, PHP, and Python) and up to 42 programming language pairs, which was collected from the GeeksForGeeks website (gee, [n. d.]). The platform ensures its contributors stick to a template in terms of the comments used in their programs and the code corresponding to those comments. The dataset provides a good number of multilingual instances of code that can be effectively used for this task. The detailed statistics of the dataset are shown in Table 1, while the train, validation, and test sets are split the same as CoST (Zhu et al., 2022). For each language, there are some instances of corresponding samples that are missing, especially in C and PHP.
@@ 4.3. RQ1: What is the Performance of VIM-PT Comparing With Baseline Approaches?
It can be observed that VIM-PT performs performance gains widely over the state-of-the-art approaches on most of the translation pairs, especially at the program level. VIM-PT performs better with the improvement of 4.57% and 1.87% over the state-of-the-art approach in multi-lingual program translation (Zhu et al., 2022) at the program level and snippet level, respectively. In particular, VIM-PT performs better than all the pairwise approaches and multi-lingual approaches on all the translations when Python is the source language or target language. Table 4 further calculates the average score of the translation of different approaches from or to any programming languages based on Table 3. As shown in Table 4, VIM-PT has performed the state-of-the-art at both snippet and program level in all configurations of the average score from or to any programming languages. We notice the performance of MusT-PT is lower than M-PT(w/o VI) on PHP and Python, which indicates that using the weight-sharing decoder is not conducive to the modeling of low-resource languages with few samples, which may be overwhelmed by the rich-resource languages in the joint training.
@@ 4.5. RQ3: Which rich-resource languages benefit the low-resource languages?
Due to the diversity between different programming languages, we explore the auxiliary benefits of different programming languages as rich-resource to the translation between low-resource programming languages. As shown in Figure 7, with auxiliary of any other programming languages, the performance of each direction of program translation has obtained a significant improvement. It is because when importing the third programming language, more data are utilized in the learning and the language-shared latent space can be constructed better. The base indicates the pair-wised approach based on CodeBERT and the others are the translation with the help of the other languages based on the framework of VIM-PT. The auxiliary benefits of different programming languages may depend on the following possible factors.
@@ 4.5.1. The imbalance of the data of different auxiliary languages
The amount of data of auxiliary language has a significant influence on the auxiliary benefit of program translation. As mentioned above, more data are utilized in the learning and the language-shared latent space can be constructed better, and further influence the decoding of the language-shared features. As illustrated in Figure 5, compared to PHP the number of pairs from Java, C++, and C# is larger. Therefore, when Java, C++, and C# are selected as auxiliary languages, the benefits for the translation are more significant than PHP as the auxiliary language generally.
@@ 4.5.2. Compiled and interpreted programming languages
Programming languages can be divided into compiled and interpreted types, depending on how programs are executed. Compiled languages require that all source code be translated into machine code when writing a program, which is then executed by a computer. Common compiled languages include C, C++, and Java. Interpreted languages translate source code into machine code line by line or block by block when executing a program and execute it. Common interpreted languages include Python, JavaScript, and PHP. As shown in Figure 7, in the compiled-compiled translation from C++ to Java, using the compiled language C# as auxiliary can achieve higher performance than the other interpreted languages, such as PHP, Javascript, and Python.
@@ 4.5.3. Syntax Similarity
Although each programming language has a unique formal syntax, the syntax similarity between different languages is various. For instance, C# is similar to C++ in syntax, while quite different from Python. As shown in Figure 7, in the translation from PHP to C++, with the help of C#, the improvement of the performance is more significant than Python as an auxiliary.
@@ 4.6. RQ4: What is the deployment complexity of the joint model and pairwise approaches?
To compare the deployment complexity between the pairwise approach and VIM-PT, we record the space overhead of the translation model. As shown in Figure 8, the pairwise approach and VIM-PT are both constructed with CodeBERT as encoder and the 6-layer Transformer as Decoder. With the number of languages in translation increasing, the number of parameters of the pairwise approach grows exponentially, and VIM-PT grows linearly.
@@ 5. Threat Analysis
  • The external threat to validity lies in the dataset used in the experiment. To mitigate the external threat, the widely-used dataset is used to evaluate the effectiveness of the method, which includes 7 programming languages and up to 42 pairs and two levels. The artifact contains the dataset and the code is publicly available in the supplementary material.
@@ 7. Acknowledgments
This research was supported by NSFC (62076121, 61921006), Major Program (JD) of Hubei Province (2023BAA024), and Postgraduate Research & Practice Innovation Program of Jiangsu Province.

### v2

@@ Abstract.
submissionid: 2705journalyear: 2024copyright: acmlicensedconference: 39th IEEE/ACM International Conference on Automated Software Engineering ; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3695553isbn: 979-8-4007-1248-7/24/10ccs: Software and its engineering Automatic programming
@@ 1. Introduction
Programs form the foundation of computer applications. Various programming languages have been invented to address diverse requirements. Developing a new application from scratch is time- and labor-intensive, whereas referring to and combining existing code is more efficient. Unfortunately, when combining programs written in different programming languages, the developer usually suffers from the intensive labor of manually translating programs from one programming language to another. For example, many industries spend hundreds of millions of dollars to convert code written in older programming languages (e.g., FORTRAN and COBOL) to newer ones (e.g., Java, C++) (Rozière et al., 2020). To alleviate the burden of program migration and facilitate the development of software systems, program translation, which aims to translate the program from one programming language to another automatically, has drawn significant attention in the software mining community (Ledley and Wilson, 1962; Kennedy III, 1987; Ahmad et al., 2023; Zhu et al., 2023; Yan et al., 2023; Weisz et al., 2022; Rozière et al., 2020; Zhu et al., 2022; Waters, 1988; Chu, 1993).
@@
In this paper, we argue that jointly training a unified model to translate code across multiple programming languages is superior to separately training with pairwise data. A program implemented in different programming languages should exhibit the invariant language-shared semantics, i.e., perform the same tasks. Thus, joint training on multiple different implementations of a program can complement and enhance the learning of the language-shared representation. By constructing the unified language-shared latent space, the rich resource languages can benefit the low resource languages in the translation. However, joint learning involves a new challenge: as shown in Fig. 1, practical datasets are always semi-parallel, which contain only a few complete multi-parallel samples that include all languages, while most are partially missing implementations in some programming languages. To tackle the issue of semi-parallel data, utilizing semi-supervised learning techniques is intuitive. Generative semi-supervised learning offers a solution: all data, including partially missing and multi-parallel data, can be generated from a latent distribution. All these data can be used to learn the latent distribution and generative process with an Expectation–Maximization (EM)-based algorithm (Kingma et al., 2014). Once the latent distribution and generative process are learned, program translation can be achieved with a conditional generative process.
@@
    We conduct extensive experiments to demonstrate the effectiveness of VIM-PT and the necessity of jointly learning a unified model in multilingual program translation.
@@ 2.1. Program Translation
To meet some specific requirements, rule-based translation methods have been developed (Alsubhi et al., 2019; An et al., 2018; Pino et al., 2017; Tauda et al., 2021). Further, some works applied phrase-based statistical machine translation techniques to program translation (Karaivanov et al., 2014; Nguyen et al., 2015, 2013, 2016; Allamanis et al., 2018; Oda et al., 2015b, c, a), which leveraged grammatical structures of programming languages for code migration. Recently, various deep learning techniques were employed to program translations (Lample et al., 2018; Wen et al., 2022; Weisz et al., 2021), which can be distinguished into pairwise program translation methods (Chen et al., 2018; Guo et al., 2021; Du et al., 2023; Yang et al., 2024) and multilingual program translation methods (Zhu et al., 2022) by using bilingual or multilingual corpus in the supervised learning.
The pairwise approaches trained one model for each translation direction independently (Guo et al., 2021; Sun et al., 2022; Liu et al., 2024). However, a problem encountered by many existing models is that program translation datasets are usually not balanced in size for all the languages. Some languages may have much less parallel data than others. Less parallel training data can significantly affect the translation performance of low-resource languages. Therefore, some works leveraged multilingual training to improve the performance of low-resource languages, which is defined as multilingual program translation. Aligning the shared semantics of multi-parallel data is the key to utilizing the rich resource to benefit the low resource. However, the existing multilingual program translation model MuST-PT (Zhu et al., 2022) is limited to the pairwise training process and does not take full advantage of the multi-parallel data and partially missing data (e.g., non-parallel data).
@@ 2.2. Exploring Non-Parallel Data for Translation
The quality of translation systems highly depends on the quality of the available parallel data. However, for most languages, parallel resources are rare or nonexistent. Some approaches have investigated the use of non-parallel data to improve existing program translation systems (Rozière et al., 2020, 2022; Szafraniec et al., 2022; Lachaux et al., 2021; Feng et al., 2020; Wang et al., 2021). For instance, Lachaux et al. (2021) proposed DOBF that leverages the structural aspect of coding languages and pre-trains a model to recover the original version of obfuscated source code. Feng et al. (2020) proposed CodeBERT, which is a Bert-like model pre-trained in an open-source GitHub repository and used in many down-stream tasks, and then much relevant research emerged (Guo et al., 2021; Du and Yu, 2023; Ma et al., 2023). Rozière et al. (2020) proposed TransCoder, which trained a translation model only using the monolingual corpus with back-translation and denoising auto-encoding.
@@ 3.1. Generative Model via Variational Inference
Consider a sample program implemented in $N$ languages, denoted as $(x^{1},x^{2},\cdots,x^{N})\sim p_{D}(x^{1},x^{2},\cdots,x^{N})$, where $x^{i}$ represents the code written in the $i$-th programming language. All of these $N$ codes exhibit the same semantics, i.e., perform the same task, while each code $x_{i}$ involves the language-specific grammatical style. From the perspective of disentanglement, each code is generated through the interplay of a language-shared feature, denoted as $z^{s}\in Z^{s}$, and a language-specific feature, denoted as $z^{i}\in Z^{i}$ for the $i$-th programming language.
@@ 4.1. Dataset Description
We conduct experiments on the CoST dataset (Zhu et al., 2022), which is a large and comprehensive dataset to evaluate the performance of program translation approaches. The dataset consists of both snippet-level and program-level parallel data from 7 programming languages (i.e., C, C#, C++, Java, JavaScript, PHP, and Python) and up to 42 programming language pairs, which was collected from the GeeksForGeeks website [^1] https://www.geeksforgeeks.org/. The platform ensures its contributors stick to a template in terms of the comments used in their programs and the code corresponding to those comments. The dataset provides a good number of multilingual instances of code that can be effectively used for this task. The detailed statistics of the dataset are shown in Table 1, while the train, validation, and test sets are split the same as CoST (Zhu et al., 2022). For each language, there are some instances of corresponding samples that are missing, especially in C and PHP.
Fig. 5 and Fig. 6 have illustrated the imbalanced proportion of the pairwise data in the dataset. The statistics of the pairwise data are shown in the Table 2 in detail.
@@
Table 5. Comparison with LLMs on program-level, including CodeT5, Deepseek-Coder(1.3B), and GPT-3.5.
Model Py-Java C#-Java C#-Py C++-C# C#-C++
CodeT5 66.40 82.41 62.80 72.47 78.39
Deepseek-Coder 54.13 52.97 59.00 55.07 53.03
GPT-3.5 65.60 85.64 62.93 79.01 81.31
VIM-PT 67.64 87.63 65.78 83.27 84.82
@@ 4.3. RQ1: What is the Performance of VIM-PT Comparing With Baseline Approaches?
It can be observed that VIM-PT performs performance gains widely over the state-of-the-art approaches on most of the translation pairs, especially at the program level. VIM-PT performs better with the improvement of 4.57% and 1.87% over the state-of-the-art approach in multi-lingual program translation (Zhu et al., 2022) at the program level and snippet level, respectively. In particular, VIM-PT performs better than all the pairwise approaches and multi-lingual approaches on all the translations when Python is the source language or target language. Table 4 further calculates the average score of the translation of different approaches from or to any programming languages based on Table 3. As shown in Table 4, VIM-PT has performed the state-of-the-art at both snippet and program level in all configurations of the average score from or to any programming languages. Additionally, the p-values between VIM-PT and each baseline are less than 0.05, confirming that the observed differences are statistically significant. We notice the performance of MusT-PT is lower than M-PT(w/o VI) on PHP and Python, which indicates that using the weight-sharing decoder is not conducive to the modeling of low-resource languages with few samples, which may be overwhelmed by the rich-resource languages in the joint training.
Moreover, we further evaluate the performance of large language models on the program translation. As shown in Table 5, CodeT5 (Wang et al., 2021) and VIM-PT are trained by supervised fine-tuning, while Deepseek-Coder(1.3B) (Guo et al., 2024) and GPT-3.5 (OpenAI, 2023) are evaluated by inference. Limited by space, we only present some of the experimental results. As shown in Table 5, GPT-3.5 performs well across language pairs with direct inference, whereas the performance of Deepseek-Coder(1.3B) is comparatively lower, which can be attributed to differences in model size. But VIM-PT achieves comparable results to GPT-3.5 despite having significantly fewer parameters.
@@ 4.5. RQ3: Which rich-resource languages benefit the low-resource languages?
Due to the diversity between different programming languages, we explore the auxiliary benefits of different programming languages as rich-resource to the translation between low-resource programming languages. As shown in Fig. 7, with auxiliary of any other programming languages, the performance of each direction of program translation has obtained a significant improvement. It is because when importing the third programming language, more data are utilized in the learning and the language-shared latent space can be constructed better. The base indicates the pair-wised approach based on CodeBERT and the others are the translation with the help of the other languages based on the framework of VIM-PT. The auxiliary benefits of different programming languages may depend on the following possible factors.
@@ 4.5.1. The imbalance of the data of different auxiliary languages
The amount of data of auxiliary language has a significant influence on the auxiliary benefit of program translation. As mentioned above, more data are utilized in the learning and the language-shared latent space can be constructed better, and further influence the decoding of the language-shared features. As illustrated in Fig. 5, compared to PHP the number of pairs from Java, C++, and C# is larger. Therefore, when Java, C++, and C# are selected as auxiliary languages, the benefits for the translation are more significant than PHP as the auxiliary language generally.
@@ 4.5.2. Compiled and interpreted programming languages
Programming languages can be divided into compiled and interpreted types, depending on how programs are executed. Compiled languages require that all source code be translated into machine code when writing a program, which is then executed by a computer. Common compiled languages include C, C++, and Java. Interpreted languages translate source code into machine code line by line or block by block when executing a program and execute it. Common interpreted languages include Python, JavaScript, and PHP. As shown in Fig. 7, in the compiled-compiled translation from C++ to Java, using the compiled language C# as auxiliary can achieve higher performance than the other interpreted languages, such as PHP, Javascript, and Python.
@@ 4.5.3. Syntax Similarity
Although each programming language has a unique formal syntax, the syntax similarity between different languages is various. For instance, C# is similar to C++ in syntax, while quite different from Python. As shown in Fig. 7, in the translation from PHP to C++, with the help of C#, the improvement of the performance is more significant than Python as an auxiliary.
@@ 4.6. RQ4: What is the deployment complexity of the joint model and pairwise approaches?
To compare the deployment complexity between the pairwise approach and VIM-PT, we record the space overhead of the translation model. As shown in Fig. 8, the pairwise approach and VIM-PT are both constructed with CodeBERT as encoder and the 6-layer Transformer as Decoder. With the number of languages in translation increasing, the number of parameters of the pairwise approach grows exponentially, and VIM-PT grows linearly.
@@ 5. Threat Analysis
  • The external threat to validity lies in the dataset used in the experiment. To mitigate the external threat, the widely used dataset, which includes 7 programming languages and up to 42 pairs and two levels, is used to evaluate the effectiveness of the method. The artifact contains the dataset and the code is publicly available in the supplementary material.
@@ 6. Conclusion
7. Data Availability
Our replication package (including code, model, etc.) is publicly available at https://github.com/duyali2000/VIM-PT.
@@ 8. Acknowledgments
This research was supported by NSFC (62076121, 61921006), Major Program (JD) of Hubei Province (2023BAA024), and Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX24_0301). The authors would like to thank Hao-Yuan He for his helpful feedback on drafts of the paper.